part 1 Find the present values of the following cash flow streams at a 10% discount rate. Do not round intermediate calculations. Round your answers to the nearest cent.

0 1 2 3 4 5
Stream A $0 $100 $350 $350 $350 $300
Stream B $0 $300 $350 $350 $350 $100

Stream A $  

Stream B $  

  1. What are the PVs of the streams at a 0% discount rate? Round your answers to the nearest dollar.

    Stream A $  

    Stream B $  

part 2

An investment will pay $150 at the end of each of the next 3 years, $200 at the end of Year 4, $300 at the end of Year 5, and $600 at the end of Year 6. If other investments of equal risk earn 5% annually, what is its present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present value: $   

Future value: $   
Answer #1

1. a . PV of Stream A =Cash Flow 1/(1+r)+Cash Flow 2/(1+r)^2+Cash Flow 3/(1+r)^3+Cash Flow 4/(1+r)^4+Cash Flow 5/(1+r)^5
=100/(1+10%)+350/(1+10%)^2+350/(1+10%)^3+350/(1+10%)^4+300/(1+10%)^5=1068.46
b. PV of Stream B =Cash Flow 1/(1+r)+Cash Flow 2/(1+r)^2+Cash Flow 3/(1+r)^3+Cash Flow 4/(1+r)^4+Cash Flow 5/(1+r)^5
=300/(1+10%)+350/(1+10%)^2+350/(1+10%)^3+350/(1+10%)^4+100/(1+10%)^5=1126.09

at 0% Rate =PV of Stream A =100+350+350+350+300 =1450
at 0% Rate =PV of Stream B =300+350+350+350+100 =1450

2. PV of Investment =150/(1+5%)+150/(1+5%)^2+150/(1+5%)^3+200/(1+5%)^4+300/(1+5%)^5+600/(1+5%)^6 =1255.8148
or 1255.81
FV of Investment =PV of Investment*(1+r)^n =1255.8148*(1+5%)^6=1682.91